Skip to content
View in the app

A better way to browse. Learn more.

Twins Daily

A full-screen app on your home screen with push notifications, badges and more.

To install this app on iOS and iPadOS
  1. Tap the Share icon in Safari
  2. Scroll the menu and tap Add to Home Screen.
  3. Tap Add in the top-right corner.
To install this app on Android
  1. Tap the 3-dot menu (⋮) in the top-right corner of the browser.
  2. Tap Add to Home screen or Install app.
  3. Confirm by tapping Install.

Game Thread Twins vs Tigers 6:40pm 9/22/2020

Featured Replies

WOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!!!!!!!!!!!!!!!!!!

  • Replies 197
  • Views 12.8k
  • Created
  • Last Reply

Top Posters In This Topic

Most Popular Posts

Should have walked him 

Whew, that was almost a disaster. Good thing the bats woke up, and Kepler seems to be getting his groove back. 

The Twins sure know how to win and give us a heart attack before doing it.

 

Now catch some breath.

Come-from-behind wins are always nice, but a come-from-behind-from-ahead win like this is maybe even a little sweeter.

 

It's almost like getting the win that goes in the books plus the win they thought they blew when they gave away the lead.

 

Either way...

 

thai-styled-bbq-chicken-taste-154759-2.j

Happy Fall everyone

Im buying a robe as soon as they are released

 

Dicknjustin were spouting the fallacy that the runner-on-second rule shortens games. The two teams have an equal chance at scoring whether there's a runner on second or not. It makes it more likely that runs will score but it gives neither team an advantage and therefore does not shorten games.

I think I disagree.   Humor me for a minute, and assume that the chance of scoring with no runner is 10% and the chance of scoring with a runner 25%.  The odds that a game goes to the 11th is one minus the odds that the away team scores and the home team doesn't  (.1X.9=.09) or the away team doesn't and the home team does (also .09.)   So the odds of going to another inning is 1-.09-.09 or 82%.   Under the new format, the odds of the away team scoring and the home team not is (.25x.75=.1875), and the odds that the away team does not and the home team does is the same.   So the odds of going to another inning is 1-.1875-.1875 or 62.5%.   I've simplified to ignore the effect of multiple run innings, but I don't think it changes the qualitative result.   The runner does speed up the game.

 

That doesn't mean it's a good change in rules.  You could speed up the game even more by flipping a coin at the end of nine, and that's a horrible idea.

 

Dicknjustin were spouting the fallacy that the runner-on-second rule shortens games. The two teams have an equal chance at scoring whether there's a runner on second or not. It makes it more likely that runs will score but it gives neither team an advantage and therefore does not shorten games.

Pretty sure it does shorten games. Huge edge for the home team if they hold the visitors without scoring in the top of the innings. I know it's only a 60 game season, but I don't recall a game going beyond 12 inning and none of those 18-20 inning marathons that happen every week or so.

 

Do you have time to feel panic or the adrenaline spike as it's on its way, or does that happen immediately after it passes/hits?

Happened after for me. 

Account

Navigation

Search

Search

Configure browser push notifications

Chrome (Android)
  1. Tap the lock icon next to the address bar.
  2. Tap Permissions → Notifications.
  3. Adjust your preference.
Chrome (Desktop)
  1. Click the padlock icon in the address bar.
  2. Select Site settings.
  3. Find Notifications and adjust your preference.